OpenStudy (firejay5):

Show work!!!!!!!!! 10. While in battle, a 15kg steel sword strikes across a horizontal steel shield. What is the normal force acting on the sword?

OpenStudy (firejay5):

My Answer so far....: Fn = m*g Fn = (15)*(9.8) Fn = 147 N

OpenStudy (firejay5):

@harishk @midhun.madhu1987

OpenStudy (midhun.madhu1987):

Yes... it is correct

OpenStudy (anonymous):

wait, you said across horizontally right.. if it has any angle mentioned .. if so, you might have to use this cos(*) thing after the formula.. propably...